the one's digit of a two digit number is one more than the ten's digit. If the digits of the number are reversed, the new number is 3 less than twice the original. Find the number.
Answer by mananth(16949)

the one's digit of a two digit number is one more than the ten's digit. If the digits of the number are reversed, the new number is 3 less than twice the original. Find the number.
let the ten's digit be x
the unit digit will be y
y=x+1
y-x=1
10y+x= 2(10x+y)-3
10y+x= 20x+2y-3
8y-19x=-3
.
8y-8x -8y+19x= 8+3
11x=11
x=1
y=x+1
y=2
the number = 12